A meter stick swinging from one end oscillates with a frequency f0. What would be the frequency, in terms of f0, if the bottom third of the stick were cut off?

Grade:11

2 Answers

Sahabudin
7 years ago
when
DR STRANGE
7 years ago
HI,
Its conservation of angular momentum concept! I1f1=I2f2............(1)
moment of inertia of rod (stick) in this case is, I= ML2/3. L=lenth of rod(stick)
I1=ML2/3 f1=f0
the bottom third of the stick were cut ,final length = 2L/3 final mass =2M/3
I2=(2M/3)(2L/3)2/3 =8ML2/81 put all this values in (1)
get f2=27f0/8 ans.